OpenStudy (anonymous):

55 mi / h = ____ ft / s Convert each rate using dimensional analysis. Round to the nearest tenth, if necessary.

OpenStudy (lucaz):

1 mile = 5280 ft 55 miles = 55*5280 1 hour = 3600 s converting: 55mi/h = 55*5280 ft / 3600 s = calculate it

OpenStudy (anonymous):

80.7

OpenStudy (lucaz):

yes..